Grades assigned by an economics lecturer for his postgraduate coursework students at a SLIIT have historically followed a symmetrical distribution: 5% HDs, 25% Ds, 40% Cs, 25% Ps and 5% Fs. This year, a sample of grades revealed 11 HDs, 32 Ds, 62 Cs, 29 Ps and 16 Fs. Can you conclude, at the 1% level of significance, that this year’s marks are distributed differently from marks in the past?


1
Expert's answer
2021-03-26T10:22:31-0400

Let "H_o" : This year’s marks are distributed differently from marks in the past.


The dstribution is as follows-


Expected frequency represented by "E_i"





So The value of "\chi^2=\sum \dfrac{(O_i-E_i)^2}{E_i}=15.12"


The tabulated value of "\chi^2" at 1% level of significance is 13.28


Conclusion: As the calculated value of "\chi^2" is greater than tabulated value So "H_o" is rejected

means This year’s marks are distributed differently from marks in the past.